what is the value of orbital angular momentum of 6s orbital ?
 

Grade:9

3 Answers

Guneet Singh
7 years ago
The value of orbital angular momentum of an electron in an orbital is independent of its principal quatum number. For any electron, the orbital angular momentum is given by the expression {l ( l + 1)} ^ 0.5 * n*h/2*pi So, for the electron in the 6s orbital, the required value is (0 * 1)^0.5 * n*h/2*pi = 0 kg m^2 s^-1
SRC Jayanth
7 years ago
The orbital angular momentum can be given as 
\frac{\displaystyle \sqrt{l(l+1)}h}{2\pi }
for s orbital l=0
by substituting orbital angular momentum = 0
 
Mridul Mohan Bagri
7 years ago
The original Angular Momentum of a orbital is given by the formula L=√l(l+1)h/2πSo for s orbital the Angular Momentum Quantum Number(l) is 0.Therefore the Orbital Angular Momentum is zero